# Translationese-index: Using Likelihood Ratios for Graded and Generalizable Measurement of Translationese

### v1

@@ 4.4 Results
We present the results of various measuring (or detecting) methods in Table 2. We mainly evaluate the cross-domain generalization, where models are trained on the single domain of Oliver Twist (OT) translated by Qwen2.5-72B-Instruct.
@@
On the in-domain test set, T-index almost perfectly classifies low-translationese and high-translationese. Two supervised methods, Bradley-Terry RM and XLM-R, also achieve an accuracy over $90\%$. When it comes to the cross-domain test sets, an increasing gap occurs between T-index and the supervised baselines. For the genre shift, T-index remains highly robust with over $95\%$ of accuracy, while a 10+ points drop is observed for Bradley-Terry RM and XLM-R, possibly due to learning domain-specific features rather than translationese features (Amponsah-Kaakyire et al., 2022).
@@
[…]
Logits might encode stylistic features in training distribution.
@@ 5 Measuring translationese in the wild
We sample 50 texts from the same source texts in the synthetic dataset and translate them into Chinese with 3 MT systems or LLMs (with the vanilla prompt) randomly selected from a pool of seven MT systems/LLMs (see Appendix B), obtaining 3 translations per source text. Following the annotation schemes in the synthetic dataset, we collect pointwise human ratings and pairwise judgments for in-the-wild MTs:
  • Pointwise: we ask 12 master students in Translation Studies in a Chinese university to rate the degree of translationese from 0-5 on 50 (source, translation) pairs. We then use the mean rating of each text as the human result.
  • Pairwise: a different set of five annotators is asked to choose the high-translationese translation, given 150 (source, translation1, translation2) triples. We take the majority vote as the ground truth.
Annotators are also asked to extract spans containing the most translationese, which can be used for token/span-level translationese measurement. Each annotation, for both pointwise and pairwise, takes about 1 min. For every 50 annotations, the annotator is compensated with 60 Chinese yuan.
@@ 5.1 T-index aligns with human judgments
For pairwise evaluation, we compute T-index for each translation and compare the value. The sample in a pair with a greater value is considered to be the prediction given by T-index. We further include several baselines that are often used in practice, including two general reward models: InternLM2-20B-Reward (Cai et al., 2024) and Skywork-Reward-Gemma2-27B-v0.2 (Liu et al., 2024); and LLM-as-a-judge methods: LLama3.3-Instruct-70B (Grattafiori et al., 2024) and Qwen2.5-Instruct-72B (Qwen et al., 2025).
@@
In the binary classification experiment, we only use data from one single domain of the synthetic dataset for SFT to test the generalization. In this section, we compare scoring models trained with different data: (1) Unpaired samples: SFT data for the two models come from two different domains; (2) Single domain: SFT data for the two models are paired, but only one domain of data is used; (3) Mixed domain: We mix pairs from all domains together with the sample size ranging from 1k to 5k.
Agreement count 3 4 5
Evaluation method #Param. N=45 N=66 N=39
T-index 2 $\times$ 0.5B
      w/ unpaired (1k) 66.6 63.6 74.3
      w/ single-dom. (1k) 66.6 69.6 82.0
      w/ mixed-dom. (1k) 64.4 71.2 87.1
      w/ mixed-dom. (3k) 66.6 72.7 87.1
      w/ mixed-dom. (5k) 62.6 78.7 89.7
InternLM2-Reward 20B 60.0 65.1 64.1
Skywork-Reward-Gemma2 27B 57.7 65.1 58.9
LLama3.3-Instruct 70B 57.7 53.0 79.4
Qwen2.5-Instruct 72B 62.2 68.1 84.6
@@
Table 3: Agreement between automated methods with humans’ majority votes in the pairwise annotation (accuracy reported). Agreement count is the number of annotators (N=5) who agree with the majority vote.
Figure 2: Correlation between the T-index and the mean human ratings for each text in the pointwise annotation. Each point represents the mean rating (0-5) on translationese from 12 annotators.
T-index agrees with human pairwise judgments.
T-index, with the pair of scoring models trained on 5k samples from mixed domains, achieves the highest accuracy of $89.7\%$ when all 5 annotators agree (see Table 3), surpassing general reward models and LLM-as-a-judge methods with up to 72B parameters. The accuracy drops for all methods as human agreement decreases, from 5 annotators agree to 4 or 3 annotators agree (see Table 3). We provide analysis on human disagreements in Appendix D.
T-index correlates with human pointwise ratings.
We also evaluate the correlation between T-index and human pointwise ratings. T-index achieves a moderate correlation with human ratings ($r=0.56$) (see Figure 2). The results suggest that the absolute scores of T-index can be used to estimate human ratings, supplementing the results in pairwise evaluation, where the relative score are able to predict human pairwise judgment.
Ablation of training data.
Ablation of training data in the upper half of Table 3 shows that the paired data is essential for T-index. Increasing the diversity of training data helps it better align with human judgment. Finally, 1-5k synthetic pairs are already sufficient, suggesting that computing T-index is data-efficient.
@@ 6 Related Work
Linguistic theories hypothesize that certain features could serve as indicators of translationese (Toury, 1995; Baker, 1996). These features are operationalized through feature engineering in the binary classification task between original and translated texts. For instance, Volansky et al. (2015) and Hu and Kübler (2021) train machine-learning classifiers with linguistic features, such as type-token ratio, POS n-grams, and grammar rules. Some information-theoretic features (Lembersky et al., 2012; Rubino et al., 2016; Bizzoni et al., 2020) are also used to identify the translationese. Therefore, these features are often used to estimate the translationese level of a text. For example, Li et al. (2025) uses perplexity, lexical density, and length variance for estiamte translationese to filter training data of higher quality. However, these methods mostly provide a corpus-level statistics, which is not applicable to MT QE that often requires sample-level information.
@@
Though nonnative, translationese it is not necessarily a defect in translation. In translation theory, translators can even purposely foreignize the translations relating to the style and culture in the source language (Venuti, 1994). For HT, translationese is not an obvious indicator of poor translation quality (Kunilovskaya and Lapshinova-Koltunski, 2019). The same goes for MT, the mild translationese is also acceptable for better faithfulness and accuracy (Freitag et al., 2022; Flamich et al., 2025). However, the rigid translationese, more frequently observed in MTs (Freitag et al., 2019; Bizzoni et al., 2020; Luo et al., 2024), is what should be penalized. Translationese is part of the overall translation quality, but it is just one of the many factors that affect the quality. However, in this paper, we isolate the translationese from other factors.
@@
7 Conclusion
In this paper, we propose T-index to quantify translationese by computing the likelihood ratios of a translation given by two contrastively fine-tuned LMs, one low-translationese model and one high-translationese model. We start with the binary classification between extremely low-translationese and high-translationese in a synthetic benchmark, where we find that T-index is able to generalize to unseen domains, since it cancels out confounding stylistic shifts caused by genre and author, making it a more generalizable measurement with translationese features isolated. We then validate T-index with human judgments in two settings: pairwise comparison and pointwise ratings. The results suggest that T-index, scored by two models of 0.5B parameters that are trained with merely 1-5k synthetic samples, can achieve around $90\%$ of agreement with human pairwise judgments when humans agree unanimously. Furthermore, there is also a moderate correlation is observed between T-indices and human pointwise ratings (pearson $r=0.568$). We then show that there is only a weak (almost none) correlation between T-index and existing QE metrics. Therefore, we conclude that T-index proposed in this paper can offer a new perspective for more holistic MT QE.
@@ 7 Conclusion
We only verified T-index on the English-Chinese language pair. Further research is needed to see whether the results are generalizable to other language pairs, especially when the two languages are similar and translationese is more difficult to define and detect. We only conduct preliminary human experiments, collecting a limited number of human annotations both in the synthetic benchmark and the in-the-wild dataset. Future research can build upon our work to collect more human annotations to examine human (dis)agreement in greater depth.
@@
The source texts used in this paper are sampled from classic novels and well-curated corpora, and the translations are produced by open-sourced LLMs. Therefore, we believe that there is no risk of leakage of personal identifiable information or any ethical issues. The data used in this paper, are only intended for research concerning the MT evaluation and should not be interpreted otherwise.
@@ Appendix A Empirical confirmation of T-index
[…]
@@
[…]
@@ B.1 Synthetic benchmark
{CJK*}
UTF8gbsn
Prompt Type Prompt Content
low-translationsee 请把以下文本翻译为中文。译文必需符合中文表达,多用小句、流水句以及中文俗语,不一定非要忠实于原文。请直接返回译文。 [Translation of the prompt: Please translate the following text into Chinese. The translation should be idiomatic Chinese, favoring shorter sentences, run-on sentences, and Chinese colloquialisms. It doesn’t have to be strictly faithful to the source text. Please return the translation directly.]
high-translationese 请把以下文本翻译为中文。译文必需忠实于原文,不要为了中文用语习惯对原文做任何修改。请直接返回译文。 [Translation of the prompt: Please translate the following text into Chinese. The translation must be faithful to the original text, and no modifications should be made to the text to fit conventions in the Chinese language. Please return the translation directly.]
Text Type Content
@@ Appendix C More details about baselines
Figure 5: Pairwise BLEU: the similarity of the two translations; $\Delta$ translationese index: the difference in translationese. The agreement count of annotators is negatively correlated with the similarity of the translation pair and positively correlated with the difference in translationese.
Appendix D Preliminary analysis of human disagreement
We ask 5 human annotators to judge which translation contains more translationese in a forced-choice manner. The majority vote is used as the ground truth. The inter-rater agreement, a Fleiss’ Kappa of 0.287, is much lower than reported in the synthetic dataset (0.840). This indicates that human annotators are less certain about the translationese in the wild. We group the human judgements into 3 categories according to the number of annotators that agree with each other (agreement count): (1) All 5 annotators agree with each other; (2) 4 annotators agree with each other; (3) 3 annotators agree with each other.
[…]
Appendix E Instruction for annotations
{CJK*}
UTF8gbsn

### v2

@@ 4.4 Results
We present the results of various measuring (or detecting) methods in Table 2. We mainly evaluate the cross-domain generalization, where models are trained on the single domain of Oliver Twist (OT) translated by Qwen2.5-72B-Instruct from English to Chinese, which is denoted as id (in-domain). Any other test set from a different genre, translated by another LLM, or from a different language pair, is denoted as ood (out-of-domain).
@@
On the in-domain test set, T-index almost perfectly classifies low-translationese and high-translationese. Three supervised methods, DPO with log-likelihood, Bradley-Terry RM and XLM-R, can also achieve an accuracy around $90\%$. When it comes to the cross-domain test sets, an increasing gap occurs between T-index and the supervised baselines. T-index remains highly discriminative under the influence of genre shift, while a drop in accuracy can be observed in supervised baselines as the genre shift away from the training domain. The limited generalizability of supervised measures is possibly due to learning domain-specific features rather than translationese features (Amponsah-Kaakyire et al., 2022). The same observation applies to the author shift. Yet in this setting, T-index also undergoes a decrease in accuracy.
When evaluated on English-Germany/French translations, the performance drops notably for even T-index. But transfer to these language pairs is non-trivial for DPO, RM, and T-index, which might result from the multilinguality of the pre-trained base model. We leave detailed cross-lingual analysis to future work.
@@
Performance varies for unsupervised methods depending on the features and models used for scoring. When $\tilde{\theta}$, fine-tuned on high-translationese, is used for scoring, logits-based methods generally outperform embedding-based methods. The simple log-likelihood can already achieve around 80% on most en-zh translations, and the other two methods also perform above random guessing. We suspect that embeddings might encode the semantics of translations, instead of lexical choice and word order that shape translationese, which can be better captured by the logits of LMs.
Logits might encode stylistic features in the training distribution.
@@ 5 Measuring translationese in the wild
5.1 Collection of human annotations
We sample 50 texts from the same source texts in the previous section and translate them into Chinese with 3 MT systems or LLMs (with the vanilla prompt) randomly selected from a pool of seven MT systems/LLMs (see Appendix B), obtaining 3 translations per source text. Note that we name these translations “in the wild" since we do not instruct the MT-systems or LLMs to produce translations of specific levels of translationese. Thus, all translations contain an arbitrary level of translationese that we need to find out, either with human annotators or T-index.
Following the annotation schemes in the synthetic dataset, we collect pointwise human ratings and pairwise judgments for in-the-wild MTs:
  • Pointwise: we ask over 30 master students in Translation Studies from prestigious Chinese universities to rate the degree of translationese from 0-5 on 150 (source, translation) pairs. Each text is rated by more than 10 raters, and then we use the mean rating of each text as the result.
  • Pairwise: a different set of five annotators is asked to choose the high-translationese translation, given 150 (source, translationA, translationB) triples. We take the majority vote as the ground truth.
Each annotation, for both pointwise and pairwise, takes 1-2 mins. For every 50 annotations, the annotator is compensated with 60 Chinese yuan.
Figure 2: Distribution of human annotation agreement patterns. Left: pointwise score differences by pairwise agreement level. Right: pairwise agreement counts by pointwise-pairwise consistency. Higher annotator consensus corresponds to larger rating differences between translation pairs.
In pairwise annotation, the inter-rater agreement (Fleiss’ Kappa = 0.287) is substantially lower than that observed in the synthetic dataset (0.840). Upon observing the annotations closely, we interpret the lower Kappa as the following: for some documents, translation A may demonstrate more translationese in certain positions, while translation B may show translationese in other places, thus making it difficult to make a decision at the document level. However, for other documents, the choice might be easier since one translation clearly demonstrate more translationese as a whole.
Therefore, we categorize pairwise judgments into three groups based on the number of agreement for a given triple: five (unanimous agreement), four, and three out of the five annotations. This categorization captures the difficulty and disagreement in pairwise annotations, which we then compare with pointwise ratings (see Figure 2).
We observe consistent trends between pairwise and pointwise annotations. As pairwise annotation certainty increases (higher agreement among annotators), the differences in pointwise ratings between the chosen and rejected translations become more pronounced. Similarly, agreement between the two annotation methods increases when the paired translations exhibit greater differences in translationese characteristics. We believe that different levels of disagreement among annotators (manifest in the pairwise agreement count) demonstrate that different shades of translationese are indeed observed by human annotators.
@@ 5.2 T-index aligns with human judgments
For pairwise evaluation, we compute T-index for each translation and compare the values. The sample in a pair with a greater value is considered to be the prediction given by T-index. We further include LLM-as-a-judge method for comparison: LLama3.3-Instruct-70B (Grattafiori et al., 2024) and Qwen2.5-Instruct-72B (Qwen et al., 2025) (see prompts in Appendix D).
@@
In §4, we only use data from one single domain of the synthetic dataset for SFT to test the generalization. In this section, we compare scoring models trained with different data: (1) Unpaired samples: SFT data for the two models come from two different domains; (2) Single domain: SFT data for the two models are paired, but only one domain of data is used; (3) Mixed domain: We mix pairs from all domains together with the sample size ranging from 1k to 5k. We also fine-tune BT RM and DPO-aligned models on the same 5k samples as T-index for comparison.
Pairw. agreement count 3 4 5 Pointw.
# samples N=45 N=66 N=39 N=150
[…]
T-index
      w/ unpaired (1k) 64.4 66.7 82.1 34.8
      w/ single-dom. (1k) 68.9 74.2 84.6 34.6
      w/ mixed-dom. (1k) 68.9 77.3 82.1 32.0
      w/ mixed-dom. (3k) 55.6 74.2 74.4 39.2
      w/ mixed-dom. (5k) 57.8 74.2 84.6 41.8
BT RM (5k) 57.8 72.7 76.9 40.7
DPO-aligned (5k) 62.2 66.7 76.9 19.7
LLama3.3-Instruct 57.7 53.0 79.4
Qwen2.5-Instruct 62.2 68.1 84.6
Human Pointwise 60.0 77.3 92.3
@@
Table 3: Agreement between automated methods with majority votes in the pairwise annotation (agreement reported) and correlation evaluated against mean ratings in pairwise annotation (Pearson’s $r$ reported).
Most automatic methods can predict human pairwise judgments.
The results in Table 3 demonstrate that most automated methods achieve above-chance performance in agreement with human pairwise judgments. Among the evaluated methods, T-index and Qwen2.5-Instruct-72B achieve the highest accuracy of 84.6% on pairs with unanimous agreement among all five human annotators. However, a performance gap of approximately 8 percentage points remains compared to the agreement between human pairwise judgments and human pointwise ratings.
T-index correlates moderately with human pointwise ratings.
When evaluated against continuous human pointwise ratings, we measure correlation strength using Pearson’s $r$. We find that T-index and BT RM achieve moderate-to-high correlations ($\sim$0.4) with human mean ratings. The results suggest that training on paired translationese data, as showcased by the synthetic datasets, can help the models capture the gradience of translationese. Notably, higher accuracy on pairwise judgments does not necessarily translate to stronger correlation with pointwise ratings. For example, T-index trained on 1k samples from a single domain achieves similar pairwise accuracy to T-index trained on 5k samples, yet lags 8 percentage points behind in pointwise correlation.
T-index is data-efficient.
Ablation experiments on training data in the upper half of Table 3 demonstrate that T-index is robust across different data conditions. Even when scoring models are trained on only 1k samples from different domains (unpaired), T-index can still effectively predict the degree of translationese. While increasing the amount of training data improves correlation with human pointwise ratings, it does not notably impact pairwise agreement performance.
@@ 6 Related Work
Linguistic theories hypothesize that certain features could serve as indicators of translationese (Toury, 1995; Baker, 1996). These features are operationalized through feature engineering in the binary classification task between original and translated texts. For instance, Volansky et al. (2015) and Hu and Kübler (2021) train machine-learning classifiers with linguistic features, such as type-token ratio, POS n-grams, and grammar rules. Some information-theoretic features (Lembersky et al., 2012; Rubino et al., 2016; Bizzoni et al., 2020) are also used to identify the translationese. Therefore, these features are often used to estimate the translationese level of a text. For example, Li et al. (2025) uses perplexity, lexical density, and length variance for estimate translationese to filter high-quality training data. However, these methods mostly provide corpus-level statistics, not applicable to MT QE which often requires sample-level information.
@@
Though nonnative, translationese it is not necessarily a defect in translation. In translation theory, translators can even purposely foreignize the translations relating to the style and culture in the source language (Venuti, 1994). For HT, translationese is not an obvious indicator of poor translation quality (Kunilovskaya and Lapshinova-Koltunski, 2019). The same goes for MT, the mild translationese is also acceptable for better faithfulness and accuracy (Freitag et al., 2022; Flamich et al., 2025). However, the rigid translationese, more frequently observed in MTs (Freitag et al., 2019; Bizzoni et al., 2020; Luo et al., 2024), is what should be penalized. Translationese is part of the overall translation quality, but it is just one of the many factors that affect the quality.
@@
7 Conclusion and future work
In this paper, we aim to develop a graded and generalizable measure of translationese. To this end, we reframe translationese measurement as a comparative task between different translations of the same source text, rather than binary classification between translated and non-translated text. Under this new formulation, among evaluated methods, T-index (likelihood ratios of two contrastively fine-tuned LLMs) has the best generalizability and alignment with human ratings and judgments.
We also show that T-index is weakly correlated with several automatic MT QE metrics, suggesting that T-index can be a complementary measure to existing MT QE metrics, which is especially important when existing MT QE metrics focus more on the accuracy and become less reliable in evaluating MTs of higher quality produced by LLMs (Agrawal et al., 2024; Kocmi et al., 2024).
Our work complements previous studies that view translationese as features distinguishing translated from non-translated texts by a new perspective: translationese can also relate to readers’ linguistic intuition directly. Building on this, future work could investigate more scalable annotation methods to capture this intuition through comprehensive human experiments. These annotations would then enable the development of finer-grained automated measures for MT system evaluation and post-training.
Beyond MT, this work can be extended to other natural language generation tasks. While binarized features like accuracy or factualness can be more easily automated, there remains a class of features that require graded measurement. Features like translationese or naturalness are more nuanced yet equally essential to the reading experience of LLM-generated texts, opening up important directions for future automated evaluation methods.
@@ 7 Conclusion and future work
We primarily verified T-index on the English-Chinese language pair. Further research is needed to see whether the results are generalizable to other language pairs, especially when the two languages are similar and translationese is more difficult to define and detect. We only conduct preliminary human experiments, collecting a limited number of human annotations both in the synthetic benchmark and the in-the-wild dataset. Future research can build upon our work to collect more human annotations to examine human (dis)agreement in greater depth.
@@
The source texts used in this paper are sampled from classic novels and well-curated corpora, and the translations are produced by open-sourced LLMs. Therefore, we believe that there is no risk of leakage of personal identifiable information or any ethical issues. The data used in this paper are only intended for research concerning the MT evaluation and should not be interpreted otherwise.
Acknowledgements
We thank our annotators for their help in data annotation. This work is funded by Shanghai Pujiang Program (22PJC063) awarded to Hai Hu and the General Program of National Natural Science Foundation of China (62176153) awarded to Rui Wang.
@@ Appendix A Empirical confirmation of T-index
Then, we obtain the MLL of the training distribution as a reference value, where $\texttt{ref\_MLL}=\mathrm{MLL}(D_{g_{i},a_{i},t_{i}};\theta_{g_{i},a_{i},t_{i}})$. With this reference, we can measure the distribution shift between $D_{g_{i},a_{i},t_{i}}$ and testing distribution $D_{g_{j},a_{j},t_{j}}$ by the difference between $\mathrm{MLL}(D_{g_{j},a_{j},t_{j};\theta_{g_{i},a_{i},t_{i}}})$ and ref_MLL.
@@
$\texttt{o\_shift}=\mathrm{MLL}(D_{g_{j},a_{j},t_{j}})-\texttt{ref\_MLL}$
$\texttt{g\_shift}=\mathrm{MLL}(D_{g_{j},a_{i},t_{i}})-\texttt{ref\_MLL}$
$\texttt{a\_shift}=\mathrm{MLL}(D_{g_{i},a_{j},t_{i}})-\texttt{ref\_MLL}$
$\texttt{t\_shift}=\mathrm{MLL}(D_{g_{i},a_{i},t_{j}})-\texttt{ref\_MLL}$
We obtain 28 models fine-tuned on the 28 datasets and evaluate them on all 28 tests, resulting in 784 observations in total. We run an OLS regression on the four types of shifts, where $\texttt{overall\_shift}\sim\texttt{genre\_shift}+\texttt{author\_shift}+\texttt{translationese\_shift}$. The linear model explains 97.3% of the variance in the overall_shift; genre_shift ($\beta=1.0129$, $p<0.001$), translationese_shift ($\beta=0.9813$, $p<0.001$), and author_shift ($\beta=0.7527$, $p<0.001$) are all significant predictors. The variance inflation factor (VIF) of the three predictors is $1.001$, $1.000$, and $1.002$, respectively, indicating that the three predictors are independent. The assumption (a) of LLR is empirically confirmed.
@@ B.1 Synthetic benchmark
Type Content
Translation Prompts
Low-translationese 请把以下文本翻译为中文。译文必需符合中文表达,多用小句、流水句以及中文俗语,不一定非要忠实于原文。请直接返回译文。
[Please translate the following text into Chinese. The translation should be idiomatic Chinese, favoring shorter sentences, run-on sentences, and Chinese colloquialisms. It doesn’t have to be strictly faithful to the source text. Please return the translation directly.]
High-translationese 请把以下文本翻译为中文。译文必需忠实于原文,不要为了中文用语习惯对原文做任何修改。请直接返回译文。
[Please translate the following text into Chinese. The translation must be faithful to the original text, and no modifications should be made to the text to fit conventions in the Chinese language. Please return the translation directly.]
Translation Examples
@@ Appendix C More details about baselines
All scoring models are trained on 1 or 2 A100-80G GPUs. It takes around 5 minutes to train a Qwen2.5-0.5B base model with the objectives including SFT, DPO, and RM. The implementations and hyperparameters for our model training can be found in our GitHub repository: https://github.com/yikang0131/TranslationeseIndex.
Appendix D Instruction for annotations and prompts used in LLM-as-a-judge
@@
Table 7: Prompts used in the LLM-as-a-judge method
【LLM-as-a-Judge: 中文翻译质量评估与解析】
任务描述: 在以下提供的两个英文句子的中文翻译版本中(A和B),请判断哪一个翻译更符合中文的表达习惯和语境。同时,请简要说明选择的理由。最后,您的回答应包含“A”或“B”,以表明您认为哪一版本更优。
示例:
英文原句:The quick brown fox jumps over the lazy dog.
翻译A:那只敏捷的棕色狐狸跳过了懒惰的狗。
翻译B:快速的棕色狐狸跃过懒散的狗。
评估解析:翻译A使用了“敏捷”和“懒惰”这两个形容词,更加形象生动,更符合中文表达习惯中的具体性和形象性。而“跃过”相较于“跳过”在中文中更具有画面感。因此,翻译A更优。
评估结果:A
实际任务:
英文原文1:{{source}}
翻译A:{{translation_A}}
翻译B:{{translation_B}}